OpenStudy (anonymous):

WILL FAN + MEDAL FOR HELP!!!! Solve. Express your answer in simplest form. x=____

OpenStudy (anonymous):

ATTACHED FILE BELOW!

OpenStudy (anonymous):

\[\frac{ 6x -1 }{ 5 } = \frac{ 3 }{ 1 }\] \[6x -1 = \frac{ 3*5 }{ 1 }\] \[6x - 1 = 15\] Find x